Folding Sheet Metal: A square is to be cut from each corner of an 8 inch by 11 inch rectangular piece of copper and the sides are to be folded up to form a box. If the area of the bottom is to be 50 inches squared, what size square should be cut from each corner?
I've tried and gotten this far:
(11-2x)(8-2x)=?
(88-22x)(-16x+4x^2)
22(4-x)-4x(4-x)
22-4x=? 4-x=?
I'm not sure what it is supposed to equal. I have tried several different answers. I have set it equal to 50 inches squared, 50 inches, 88 inches, and 44 inches. If you can help me with this problem, I would appreciate it! Thank you!

Answer by stanbon(75887)

Folding Sheet Metal: A square is to be cut from each corner of an 8 inch by 11 inch rectangular piece of copper and the sides are to be folded up to form a box. If the area of the bottom is to be 50 inches squared, what size square should be cut from each corner?
I've tried and gotten this far:
(11-2x)(8-2x)=50
(88-22x-16x+4x^2) = 50
4x^2-38x+38 = 0
2x^2 - 19x +19 = 0
x = [19 +- sqrt(19^2-4*2*19)]/4
x = [19 +- sqrt(209)]/4
x = 8.364 inches or x = 1.1358 inches
-----------
Only x = 1.1358 is realistic.
============
Cheers,
Stan H.